# Semantic-Aware Spectrum Sharing in Internet of Vehicles Based on Deep Reinforcement Learning

### v3

@@ Abstract
This work aims to investigate semantic communication in high-speed mobile Internet of vehicles (IoV) environments, with a focus on the spectrum sharing between vehicle-to-vehicle (V2V) and vehicle-to-infrastructure (V2I) communications. We specifically address spectrum scarcity and network traffic and then propose a semantic-aware spectrum sharing algorithm (SSS) based on the deep reinforcement learning (DRL) soft actor-critic (SAC) approach. Firstly, we delve into the extraction of semantic information. Secondly, we redefine metrics for semantic information in V2V and V2I spectrum sharing in IoV environments, introducing high-speed semantic spectrum efficiency (HSSE) and semantic transmission rate (HSR). Finally, we employ the SAC algorithm for decision optimization in V2V and V2I spectrum sharing based on semantic information. This optimization encompasses the optimal link of V2V and V2I sharing strategies, the transmission power for vehicles sending semantic information and the length of transmitted semantic symbols, aiming at maximizing HSSE of V2I and enhancing success rate of effective semantic information transmission (SRS) of V2V. Experimental results demonstrate that the SSS algorithm outperforms other baseline algorithms, including other traditional-communication-based spectrum sharing algorithms and spectrum sharing algorithm using other reinforcement learning approaches. The SSS algorithm exhibits a 15% increase in HSSE and approximately a 7% increase in SRS.
@@ I-A Background
In recent years, researchers have delved into Internet of vehicles (IoV) technology [1, 2, 3, 4]. As IoV gains attention, communications between vehicles and vehicles (V2V) as well as between vehicles and infrastructure (V2I) become crucial for supporting intelligent transportation systems (ITS) [5, 6, 7, 8, 9]. However, the success of IoV depends not only on the rapid development of technology but also on effective spectrum sharing.
@@
The scarcity of spectrum resources has triggered spectrum sharing issues in IoV systems [10, 11, 12], where V2V communications need to share spectrum resources with V2I communications [13, 14, 15, 16]. When vehicles engage in V2V and V2I communications simultaneously within the same limited spectrum bandwidth, they need to share the same spectrum resources [17, 18, 19]. This can result in conflicts due to competition for resources. Additionally, the dynamic and heterogeneous nature of IoV environments means that vehicles entering and leaving communication ranges may interfere with ongoing communications or access the same spectrum resources simultaneously [20], leading to conflicts and a potential decline in communication performance [21, 22, 23]. The aforementioned challenges can further impact the reliability and efficiency of communication in IoV systems.
@@
The diversity of IoV applications and the explosive growth of content pose challenges to spectrum sharing based on traditional communication [24]. This growth implies that traditional communication faces new challenges, its maximum rate is upper bounded by the Shannon limit (the theoretical limit of information transmission) [25]. The exchange of information between vehicles becomes increasingly rich and complex, especially in the high-speed mobile IoV environment [26], which surpasses the capabilities of communication methods based solely on bits [27]. This may result in a decrease in communication quality, thereby affecting the performance of IoV systems.
To address the problem of spectrum sharing based on traditional communication, the communication field has witnessed a shift from traditional communication paradigms to semantic communication. Semantic communication places the meaning of the transmitted source at its core, rather than just transmitting bits or data [28]. Because semantic communication can convey richer information beyond just bit data [29], adapt to dynamic environments [30], and reduce spectrum waste [31], this emerging communication paradigm is considered a potential solution to reduce network traffic and alleviate spectrum scarcity, especially in the context of spectrum sharing in V2V and V2I communication.
However, in the high-speed mobile IoV environment, semantic communication still faces specific challenges [32]. Firstly, determining what should be considered as semantic information and how to extract this information from different data streams is a complex problem. Secondly, communication between vehicles is typically based on bits derived from statistical knowledge, rather than semantic information from the source. Therefore, traditional bit-based communication methods are fundamentally different from semantic communication. It requires a redefinition of V2V and V2I spectrum sharing methods suitable for semantic communication. Lastly, vehicles need to make decisions based on semantic data, including the selection and allocation of spectrum resources. Due to the richer and more complex information it conveys, along with the introduction of additional decision variables and nonlinear information relationships, traditional optimization decision methods are not directly applicable to the semantic communication scenario. Thus, it is necessary to use a deep reinforcement learning (DRL) [33] decision algorithm to fully utilize semantic information.
@@
The objective of this paper is to explore the application of semantic communication in IoV, particularly in addressing spectrum sharing issues. We design a semantic-aware spectrum sharing algorithm (SSS) based on the DRL soft actor-critic (SAC) approach to maximize HSSE of V2I and enhance the success rate of effective semantic information transmission (SRS) of V2V links  [^1] The source code has been released at: https://github.com/qiongwu86/Semantic-Aware-Spectrum-Sharing-in-Internet-of-Vehicles-Based-on-Deep-Reinforcement-Learning. We will investigate the extraction of semantic information, redefine spectrum sharing, and optimize decision-making based on semantic data using DRL to enhance the performance of V2V and V2I communications.
@@ II System Model
As shown in Fig. 1, we consider a cellular-based vehicular communication network in urban areas, where $n$ vehicles are driving in the communication range of a BS. The BS is deployed in the center of the intersection. The initial positions of vehicles are determined by a spatial Poisson process, and the initial driving direction of each vehicle are set randomly. Each vehicle keeps driving with the initial direction and maintains a constant speed of $v$ km/s. When a vehicle arrives at the intersection, it has a ${p_{l}}$chance to the left direction, ${p_{r}}$chance to the right direction and ${p_{s}}$chance to the straight direction, where ${p_{l}}\%+{p_{r}}\%+{p_{r}}\%=1$. For keeping the number of vehicles to be the same in the considered scenarios, we also assume that when a vehicle crosses the environment boundary, its position is adjusted to the opposite edge. The position of each vehicle is updated every $m$ milliseconds. The BS and each vehicle are equipped with a transmitter and a receiver.
@@
At the beginning, the BS collects textual data from vehicles and pre-trains a DeepSC model based on the collected data, then broadcasts the DeepSC model to all vehicles in the considered scenario. Each vehicle generates high-data-rate entertainment service and reliable periodic safety information in the form of textual data and transforms the information to semantic information at the transmitter including semantic and channel encoding based on the DeepSC model, and then the vehicle’s transmitter transmits the semantic information to BS through a V2I link or another vehicle through a V2V link. Next, BS or another vehicle recovers the received semantic information to original textual data including semantic and channel decoding. Fig. 2 illustrates the transmission process of semantic information.
@@ II-A DeepSC Transceivers
The transmitting vehicle generates a sentence $S\left[w\right]$ with ${l_{q}}$ words in the $w$-th V2I link as shown in Fig. 1. $S\left[w\right]=\left[{{s_{1}}\left[w\right],{s_{2}}\left[w\right],\ldots,{s_{l}}\left[w\right],\ldots{s_{{l_{q}}}}\left[w\right]}\right]$, where ${{s_{l}}\left[w\right]}$ represents the $l$-th word.
@@ II-A1 Transmitter Model
All the sentences across all links are $S\left[w\right]$. The sentences are then input into the encoding part of the DeepSC transmitter in the vehicle in order to extract semantic information $X\left[w\right]$ from $S\left[w\right]$. The formula is given by:
$X\left[w\right]=c{h_{\beta}}\left({s{e_{\alpha}}{\rm{}}\left({S\left[w\right]}\right)}\right),$ (1)
where $s{e_{\alpha}}{\rm{}}\left(\cdot\right)$ and $c{h_{\beta}}\left(\cdot\right)$ are the semantic and channel encoder networks with parameters $\alpha$ and $\beta$ respectively. The semantic symbol vector is $x\left[w\right]={x_{1}}\left[w\right],{x_{2}}\left[w\right],\ldots,{x_{l}}\left[w\right],\ldots{x_{u_{q}l}}\left[w\right]$, where ${u_{q}}$ represents the average number of semantic symbols used for each word.
@@ II-A2 Transmission Model
${h_{q}}\left[w\right]=s{s_{q}}\left[w\right]\times l{s_{q}},$ (2)
@@
$SINR_{w}^{VI}\left[w\right]=\frac{{P_{w}^{VI}{h_{w,B}}\left[w\right]}}{{{\sigma^{2}}+\sum\limits_{q}{{b_{q}}\left[w\right]P_{q}^{VV}\left[w\right]{h_{q,B}}\left[w\right]}}},$ (3)
$SINR_{q}^{VV}\left[w\right]=\frac{{P_{q}^{VV}\left[w\right]{h_{q}}\left[w\right]}}{{{\sigma^{2}}+I\left[w\right]}},$ (4)
where $I\left[w\right]=P_{w}^{VI}{h_{w,q}}\left[w\right]+\sum\limits_{{q^{\prime}}\neq q}{{b_{{q^{\prime}},w}}\left[w\right]P_{{q^{\prime}}}^{VV}\left[w\right]{h_{{q^{\prime}},q}}\left[w\right]}$. ${P_{w}^{VI}}$ and ${P_{q}^{VV}}$ are transmit powers of the $w$-th V2I transmitter and the $q$-th V2V transmitter respectively over the $w$-th sub-band. ${b_{q}\left[w\right]}$ is binary spectrum allocation indicators. ${b_{q}\left[w\right]}=1$ represents that the $q$-th V2V link shares the $w$-th V2I link. Conversely, ${b_{q}\left[w\right]}=0$. The received semantic signal is represented as:
$Y\left[w\right]=H\left[w\right]X\left[w\right]+N,$ (5)
@@ II-A3 Receiver Model
The receiver includes a channel decoder and a semantic decoder, which recover the received semantic symbols. So the decoded signal can be represented as:
$\hat{S}\left[w\right]=se_{\mu}^{-1}\left({ch_{\nu}^{-1}{\rm{}}\left({Y\left[w\right]}\right)}\right),$ (6)
@@
$\begin{array}[]{l}{L_{CE}}(S\left[w\right],\hat{S}\left[w\right];\alpha,\beta,\mu,\nu)=\sum\limits_{l}{q({s_{l}}\left[w\right])}log({s_{l}}\left[w\right])\\ +(1-q({s_{l}}\left[w\right]))log(1-p({s_{l}}\left[w\right])),\end{array}$ (7)
@@
$B\left(\cdot\right)$ represents the bidirectional encoder representation of the Transformers (BERT) model [56] for sentences, where $0\leq\xi\leq 1$, $\xi=1$ indicates the highest similarity between the two sentences, $\xi=0$ implies no similarity between the two sentences.
@@ II-B Novel Metrics for Spectrum Sharing in High-Speed Mobile Vehicular Networks
where ${u_{q}}$ is the average number of semantic symbols used for each word, its unit is $semantic$ $unit$ (sut) [57], thus ${u_{q}}L$ is the semantic symbols of each sentence. $\frac{I}{{{u_{q}}L}}$ represents the semantic information contained in each semantic symbol. $W$ is equivalent to the quantity of semantic symbols transmitted per second numerically. Since the way utilizing the channel bandwidth efficiently is that the bandwidth occupied by the number of semantic symbols transmitted per second is equal to the bandwidth. ${\xi}$ represents the semantic similarity which depends on the number of semantic symbols ${u_{q}}$ and the SINR. The unit of HSR is $suts/s$. we obtain a function describing ${\xi}$ in terms of ${u_{q}}$ and SINR by fitting and denoted as $\xi=\Psi\left({{u_{q}},SINR\left[w\right]}\right)$ [57].
We observe that semantic similarity gradually increases with an increase in the number of semantic symbols and an improvement in signal-to-noise ratio. On the one hand, using words that encompass more detailed semantic information, making encoding and decoding more effective and resulting in greater semantic similarity. On the other hand, with higher SINR, semantic information will have less distortion when transmitted over wireless channels making the original semantic content better preserved during encoding and decoding, leading to increased semantic similarity. Therefore, the HSSE is expressed as:
@@
Its unit is $suts/s/Hz$. This metric quantifies the efficiency of transmitting semantic information, measured in symbols, per unit of available bandwidth. It provides insights into how effectively the communication system utilizes the available frequency spectrum for semantic information transmission. The value of $\frac{I}{L}$ depends on the type of source [57]. For the source in this paper, it is a constant that does not affect resource optimization. Therefore, this term $\frac{I}{L}$ can be omitted during optimization.

### v4

@@ Abstract
This paper investigates semantic communication in high-speed mobile Internet of Vehicles (IoV), focusing on spectrum sharing between vehicle-to-vehicle (V2V) and vehicle-to-infrastructure (V2I) communications. We propose a semantic-aware spectrum sharing (SSS) algorithm using deep reinforcement learning (DRL) with a soft actor-critic (SAC) approach. We start with semantic information extraction, redefining metrics for V2V and V2I spectrum sharing in IoV environments, introducing high-speed semantic spectrum efficiency (HSSE) and semantic transmission rate (HSR). We then apply SAC algorithm to optimize decisions V2V and V2I spectrum sharing decisions on semantic information. This optimization aims to maximize HSSE and enhance the success rate of effective semantic information transmission (SRS), including determining the optimal V2V and V2I sharing strategies, transmission power, and the length of transmitted semantic symbols. Experimental results show that the SSS algorithm outperforms other baseline algorithms, including other traditional-communication-based spectrum sharing algorithms and spectrum sharing algorithm using other reinforcement learning approaches. The SSS algorithm exhibits a 15% increase in HSSE and approximately a 7% increase in SRS.
@@ I-A Background
In recent years, researchers have delved into Internet of vehicles (IoV) technology [1, 2]. As IoV gains attention, cellular vehicle-to-everything (C-V2X) including communications between vehicles and vehicles (V2V) as well as between vehicles and infrastructure (V2I) become crucial for supporting intelligent transportation systems (ITS) [3]. C-V2X enhances security by leveraging cellular network encryption and authentication protocols, reducing side-channel attack risks and ensure secure data transmission with the rise of security issues [4]. However, the success of IoV depends not only on the rapid development of technology but also on effective spectrum sharing [5, 6].
@@
The scarcity of spectrum resources has triggered spectrum sharing issues in IoV systems [7], where V2V communications need to share spectrum resources with V2I communications [8]. The available radio spectrum is limited and must be divided among various services and users. With the rapid increase in the number of connected vehicles, the competition for this finite resource intensifies [9, 10]. Vehicles engaged in both V2V and V2I communications simultaneously within the same limited spectrum bandwidth, they need to share the same spectrum resources [11]. This can result in conflicts due to competition for resources as multiple vehicles vie for the same spectrum, potentially causing interference and degradation in communication performance [12]. Additionally, as vehicles move in and out of communication ranges, they may disrupt ongoing communications or access the same spectrum resources concurrently, further exacerbating the issue and impacting the reliability and efficiency of IoV systems.
@@
Network traffic has become increasingly critical in IoV due to the diverse applications and explosive growth of content, which challenge traditional communication methods for spectrum sharing [13, 14]. Vehicles frequently change positions and communication links, leading to fluctuating network loads. Additionally, V2V and V2I applications require timely and accurate data exchange. Network traffic congestion can result in increased latency and message loss. These issues highlight that traditional communication methods, constrained by the Shannon limit (the theoretical maximum rate of information transmission) [15], are inadequate for the IoV environment. As the exchange of information between vehicles becomes increasingly rich and complex, especially in the high-speed mobile IoV environment, which surpasses the capabilities of communication methods based solely on bits [16]. The high latency, limited bandwidth, and interference in vehicular networks can severely impact the communication system, thereby affecting the performance of IoV systems [17].
The challenge of spectrum scarcity and high network traffic is critical in IoV environments due to the large number of vehicles competing for limited spectrum resources. Semantic communication improves efficiency by focusing on contextually relevant and essential information, which minimizes unnecessary data transmission and optimizes spectrum usage [18]. This approach is particularly beneficial in dynamic vehicular networks, where conditions such as vehicle speed and traffic density change rapidly. By understanding and prioritizing the content and context of the data, semantic communication allows for real-time adjustments in resource allocation, reducing both spectrum occupancy and network traffic [19]. Additionally, traditional spectrum sharing often wastes bandwidth, while semantic communication reduces spectrum occupancy by focusing on essential data. Finally, by understanding data content and context, semantic communication enhances decision-making in spectrum-sharing algorithms through dynamic transmission power adjustments, optimal link selection, and scheduling based on data relevance and urgency [20]. This emerging communication paradigm is considered a potential solution to reduce network traffic and alleviate spectrum scarcity, especially in the context of spectrum sharing in V2V and V2I communication. Semantic information enhances security by ensuring data is only meaningful to authorized users who share the same knowledge base. This approach reduces the risk of interception and tampering, as encrypted semantic data cannot be easily interpreted without the correct context.
However, in the high-speed mobile IoV environment, semantic communication still faces specific challenges. First, defining and extracting semantic information from various data streams is complex. Second, traditional vehicle communication relies on bit-based methods derived from statistical knowledge rather than semantic data, necessitating a redefinition of V2V and V2I spectrum sharing for semantic communication. Finally, vehicles must make decisions based on semantic data for spectrum resource allocation. The complexity and richness of semantic information, combined with additional decision variables and nonlinear relationships, render traditional optimization methods inadequate. Therefore, a deep reinforcement learning (DRL) algorithm is needed to effectively leverage semantic information [21]. Among DRL methods, Deep Q-Network (DQN) is limited to discrete actions, making it unsuitable for IoV scenarios. While deep deterministic policy gradient (DDPG) can handle continuous actions, it lacks good exploration and stability. The Soft Actor-Critic (SAC) approach not only handles continuous action spaces effectively but also improves sample efficiency and training stability through entropy regularization. Additionally, SAC’s off-policy nature enables efficient learning by reusing past experiences, which is beneficial in dynamic IoV scenarios.
@@
The objective of this paper is to explore the application of semantic communication in IoV, particularly in addressing spectrum sharing issues. We design a semantic-aware spectrum sharing algorithm (SSS) based on the DRL SAC approach to maximize high-speed semantic spectrum efficiency (HSSE) of V2I and enhance the success rate of effective semantic information transmission (SRS) of V2V links [^1] The source code has been released at: https://github.com/qiongwu86/Semantic-Aware-Spectrum-Sharing-in-Internet-of-Vehicles-Based-on-Deep-Reinforcement-Learning. We will investigate the extraction of semantic information, redefine spectrum sharing, and optimize decision-making based on semantic data using DRL to enhance the performance of V2V and V2I communications.
@@ II System Model
As shown in Fig. 1, the DeepSC model enhances communication efficiency by focusing on the meaning of the transmitted data rather than just the raw data itself [35]. The model includes an encoding part (semantic and channel encoding), a decoding part (semantic and channel decoding), and a wireless channel, all sharing a common semantic knowledge base. Initially, each vehicle collects raw textual data over a wireless channel and transmits it to the base station (BS) within its range. The BS then forwards this data to a central cloud server via a high-speed optical fiber connection, ensuring minimal latency and negligible data loss. The cloud server preprocesses the data and uses it to pre-train the DeepSC model. After pre-training, the cloud server sends the DeepSC model back to the BS, which then broadcasts it to all vehicles, ensuring they have the latest model for local inference. Vehicles use the model to extract semantic information from the textual data they generate and can upload new data or model parameters back to the BS, which forwards them to the cloud server for further training and updates. This feedback loop keeps the model robust and adaptable to changing network conditions.
Figure 1: Semantic pre-training framework
As shown in Fig. 2, we consider a cellular-based vehicular communication network in urban areas, where $n$ vehicles are driving in the communication range of a BS. The BS is deployed in the center of the intersection. Both BS and vehicles are equipped with DeepSC model. The initial positions of vehicles are determined by a spatial Poisson process, and the initial driving direction of each vehicle are set randomly. Each vehicle keeps driving with the initial direction and maintains a constant speed of $v$ km/s. When a vehicle arrives at the intersection, it has a ${p_{l}}$chance to the left direction, ${p_{r}}$chance to the right direction and ${p_{s}}$chance to the straight direction, where ${p_{l}}\%+{p_{r}}\%+{p_{s}}\%=1$. For keeping the number of vehicles to be the same in the considered scenarios, we also assume that when a vehicle crosses the environment boundary, its position is adjusted to the opposite edge.
@@ II-A DeepSC Transceivers
The transmitting vehicle generates a sentence $S\left[w\right]$ with ${l}$ words in the $w$-th V2I link as shown in Fig. 2, where $S\left[w\right]=\left[{{s_{1}}\left[w\right],{s_{2}}\left[w\right],\ldots,{s_{l}}\left[w\right]}\right]$.
@@ II-A1 Transmitter Model
All the sentences across all links are $S\left[w\right]$. The sentences are then input into the encoding part of the DeepSC transmitter in the vehicle in order to extract semantic information $X\left[w\right]$ from $S\left[w\right]$. The semantic encoding part utilizes advanced NLP technique (Transformer-based encoding) to extract semantic information. This information is then encoded into a suitable format for transmission by the channel encoding part. The formula is given by:
$\vspace*{-0.2\baselineskip}X\left[w\right]=c{h_{\beta}}\left({s{e_{\alpha}}{\rm{}}\left({S\left[w\right]}\right)}\right),$ (1)
where $s{e_{\alpha}}{\rm{}}\left(\cdot\right)$ and $c{h_{\beta}}\left(\cdot\right)$ are the semantic and channel encoder networks with parameters $\alpha$ and $\beta$ respectively. The semantic symbol vector is $X\left[w\right]={x_{1}}\left[w\right],{x_{2}}\left[w\right],\ldots,{x_{l}}\left[w\right],\ldots{x_{u_{q}l}}\left[w\right]$, where $X\left[w\right]\in{R^{{u_{q}}l}}$ and ${u_{q}}$ represents the average number of semantic symbols used for each word. The vector dimension is determined by the semantic and channel encoders’ neural network. The value of $l$ is fixed based on the European parliament dataset [42], which contains predefined English sentences used for training, thus $l$ is known. Consequently, ${u_{q}}$ can be computed by dividing the overall vector of ${X\left[w\right]}$ by the corresponding sentence length $l$ from the dataset.
@@ II-A2 Transmission Model
$\vspace*{-0.2\baselineskip}{h_{q}}\left[w\right]=s{s_{q}}\left[w\right]\times l{s_{q}},$ (2)
@@
$\vspace*{-0.2\baselineskip}SINR_{w}^{VI}\left[w\right]=\frac{{P_{w}^{VI}{h_{w,B}}\left[w\right]}}{{{\sigma^{2}}+\sum\limits_{q}{{b_{q}}\left[w\right]P_{q}^{VV}\left[w\right]{h_{q,B}}\left[w\right]}}},$ (3)
$\vspace*{-0.2\baselineskip}SINR_{q}^{VV}\left[w\right]=\frac{{P_{q}^{VV}\left[w\right]{h_{q}}\left[w\right]}}{{{\sigma^{2}}+I\left[w\right]}},$ (4)
where $I\left[w\right]=P_{w}^{VI}{h_{w,q}}\left[w\right]+\sum\limits_{{q^{\prime}}\neq q}{{b_{{q^{\prime}},w}}\left[w\right]P_{{q^{\prime}}}^{VV}\left[w\right]{h_{{q^{\prime}},q}}\left[w\right]}$. ${P_{w}^{VI}}$ and ${P_{q}^{VV}}$ are transmit powers of the $w$-th V2I transmitter and the $q$-th V2V transmitter respectively over the $w$-th sub-band. ${b_{q}\left[w\right]}$ is binary spectrum allocation indicators. ${b_{q}\left[w\right]}=1$ represents that the $q$-th V2V link shares the $w$-th V2I link. Conversely, ${b_{q}\left[w\right]}=0$. The encoded semantic information, suitable for transmission, is sent from the vehicle’s transmitter either to BS) via V2I link or directly to another vehicle via V2V link. The received semantic signal is represented as:
$\vspace*{-0.2\baselineskip}Y\left[w\right]=H\left[w\right]X\left[w\right]+N,$ (5)
@@ II-A3 Receiver Model
The receiver includes a channel decoder and a semantic decoder. Upon receiving the encoded semantic information, the BS or the receiving vehicle uses the channel decoding and semantic decoding parts to decode it back into its original textual form. So the decoded signal can be represented as:
$\vspace*{-0.2\baselineskip}\hat{S}\left[w\right]=se_{\mu}^{-1}\left({ch_{\nu}^{-1}{\rm{}}\left({Y\left[w\right]}\right)}\right),$ (6)
@@
$\vspace*{-0.3\baselineskip}\begin{array}[]{l}{L_{CE}}(S\left[w\right],\hat{S}\left[w\right];\alpha,\beta,\mu,\nu)=\sum\limits_{l}{q({s_{l}}\left[w\right])}log({s_{l}}\left[w\right])\\ +(1-q({s_{l}}\left[w\right]))log(1-p({s_{l}}\left[w\right])),\end{array}$ (7)
@@
$B\left(\cdot\right)$ represents the bidirectional encoder representation of the Transformers (BERT) model [45] for sentences, where $0\leq\xi\leq 1$, $\xi=1$ indicates the highest similarity between the two sentences, $\xi=0$ implies no similarity between the two sentences. $\xi$ measures how closely two sentences convey the same meaning, regardless of their wording. For instance, if ”The oranges are sweet and delicious” is received as ”Sweet oranges are very fragrant”, traditional methods might see this as a failure due to wording differences. However, semantic similarity assesses if the underlying meaning is preserved, indicating successful transmission if $\xi$ is high. This metric is crucial for evaluating semantic communication performance, reflecting how well the intended meaning is maintained.
@@ II-B Novel Metrics for Spectrum Sharing in High-Speed Mobile Vehicular Networks
where ${u_{q}}$ is the average number of semantic symbols used for each word, its unit is semantic unit (sut) [46], thus ${u_{q}}L$ is the semantic symbols of each sentence. $\frac{I}{{{u_{q}}L}}$ indicates the semantic information contained in each semantic symbol. $W$ corresponds to the number of semantic symbols transmitted per second. Efficient utilization of the channel bandwidth means that the bandwidth occupied by the transmitted semantic symbols per second equals the available bandwidth. According to [46], the function is defined as $\xi=\Psi\left({{u_{q}},SINR\left[w\right]}\right)$, meaning that semantic similarity depends on the number of semantic symbols ${u_{q}}$ and SINR. In the high-mobility vehicular environment, we can calculate the semantic similarity based on the current SINR and the chosen ${u_{q}}$. The value of ${\xi}$ increases with more ${u_{q}}$ and better $SINR$. Detailed semantic info and higher $SINR$ enhance similarity by reducing distortion.
The unit of HSR is $suts/s$. HSR measures the rate at which semantic information is successfully transmitted over the wireless channel, reflecting the quality of communication. In high-speed vehicular environments, maintaining high communication quality is essential for applications such as safety alerts and real-time navigation. HSR ensures that critical information is transmitted accurately and quickly, thereby enhancing the reliability and effectiveness of the communication system.
Therefore, the HSSE is expressed as:
@@
Its unit is $suts/s/Hz$. This metric quantifies the efficiency of transmitting semantic information, measured in symbols, per unit of available bandwidth. It provides insights into how effectively the communication system utilizes the available frequency spectrum for semantic information transmission. HSSE is related to semantic similarity ${\xi}$ according to (32), and ${\xi}$ is highly correlated with the SINR [46]. Since SINR measures the power ratio between signal, noise, and interference, it is directly relevant to energy consumption. Therefore, HSSE provides a holistic view of both semantic communication efficiency and energy consumption. This metric is vital for ensuring that the communication system makes the best use of the limited spectrum resources, which is particularly important in the context of IoV where spectrum scarcity is a common issue. By focusing on the semantic content of the transmitted data, HSSE helps to maximize the information density within the available spectrum, leading to more efficient spectrum utilization. The value of $\frac{I}{L}$ depends on the type of source [46]. For the source in this paper, it is a constant that does not affect resource optimization. Therefore, this term $\frac{I}{L}$ can be omitted during optimization.